# Adaptive Knowledge Graphs Enhance Medical Question Answering: Bridging the Gap Between LLMs and Evolving Medical Knowledge

### v1

@@ 4.2 Performance Comparison
Table 1 presents a comprehensive comparison of state-of-the-art language models on the MEDQA benchmark. The results highlight the critical role of advanced reasoning strategies in achieving higher performance, such as CoT reasoning, fine-tuning, and the integration of search tools. While larger models like Med-Gemini and GPT-4 achieve the highest accuracy and F1 scores, their performance comes at the cost of significantly larger parameter sizes. These models exemplify the power of scaling combined with sophisticated reasoning and retrieval techniques.
Significantly, AMG-RAG, despite having just 8 billion parameters, attains an F1 score of 74.1% on the MEDQA benchmark, surpassing models like Meditron, which possess 70 billion parameters without needing any fine tuning. This highlights AMG-RAG’s exceptional efficiency and proficiency in utilizing CoT reasoning and external evidence retrieval. The model leverages tools such as PubMedSearch and WikiSearch to dynamically integrate domain-specific knowledge dynamically, thereby improving its ability to address medical questions. Examples of QA interactions, including detailed search items and reasoning for question samples, are provided in Appendix C. These examples are organized in Tables 6, 7, 8, and 9, drawn from the MEDQA benchmark.
@@
Overall, AMG-RAG’s results on MEDQA and MedMCQA benchmarks solidify its position as a highly efficient and effective model for medical QA. By leveraging CoT reasoning, search tools, and external knowledge sources, AMG-RAG not only closes the gap with much larger models but also sets a new standard for performance among smaller-sized models.
4.3 Impact of Search Tools and CoT Reasoning on AMG-RAG Performance
@@
Figure 3 and Table 3 demonstrate the impact of integrating search tools such as PubMedSearch and WikiSearch on the performance of AMG-RAG when applied to the MEDQA dataset. The inclusion of these search capabilities significantly improves accuracy and F1 scores by providing access to relevant external evidence, which is critical for addressing medical questions. Among the search tools, PubMedSearch outperforms WikiSearch, likely due to its more focused and domain-specific content, which better aligns with the nature of medical QA tasks.
Additionally, the impact of CoT reasoning and MKG integration on AMG-RAG performance is highlighted in the same figure and table. The results reveal that the removal of either CoT reasoning or KG integration leads to a substantial drop in accuracy and F1 scores. This underscores the indispensable role of structured reasoning and domain-specific retrieval in enhancing the system’s ability to generate accurate and evidence-backed answers.
Table 3: Performance metrics for AMG-RAG model with and without CoT and Knowledge Graph integration with different search tools for MEDQA dataset.
Model Acc. (%) F1-Score Recall
PubMedSearch 73.92 0.7410 0.7392
WikiSearch 70.62 0.7067 0.7062
No Search 67.16 0.6696 0.6716
No Search & CoT 66.69 0.6655 0.6669
@@
Figure 3: Confusion matrix for AMG-RAG with and without CoT and Knowledge Graph integration on MEDQA dataset.
4.4 Improving QA in Rapidly Changing Medical Domains
Figure 4 illustrates the performance of various models across different question domains, including Neurology and Genetics. The AMG-RAG model consistently outperforms other approaches, showcasing its superior adaptability and robustness in these rapidly evolving and knowledge-intensive fields. This exceptional performance stems from its ability to seamlessly integrate external sources of information and evidence. By leveraging PubMed searches, the AMG-RAG model dynamically retrieves the latest medical research and continuously updates the MKG, ensuring that it remains relevant and up-to-date. This dynamic updating process not only enhances the model’s ability to reason across multiple domains but also allows it to address complex, multi-hop questions with greater accuracy and depth.
@@ 5 Conclusion
In this work, we introduce AMG-RAG, an advanced QA system that dynamically constructs MKG while integrating sophisticated reasoning and external domain-specific search tools. The model exhibits significant improvements in accuracy and reasoning capabilities, particularly for medical question-answering tasks, outperforming other approaches of similar model size or 10 to 100 times larger. Using structured knowledge representations and advanced reasoning frameworks, our approach establishes a new benchmark for QA in highly competitive and highly evolving domains such as medicine.
@@ 6 Limitations
Despite AMG-RAG advancements, our approach has certain limitations. Firstly it relies on external search tools to introduce latency during the creation of MKG. However, this occurs only once, when the MKG is built from scratch for the first time. Additionally, while the model performs exceptionally well in medical domains, its applicability to non-medical tasks remains unexplored.
@@ Appendix B Evaluating the Accuracy and Robustness of the Medical Knowledge Graph
To assess accuracy and robustness, a two-phase evaluation process was employed. In the first phase, a group of expert LLMs specializing in medical domains reviewed a subset of the MKG, including dynamically generated nodes, relationships, confidence scores, and summaries for various medical queries. They evaluated the accuracy of medical terms and concepts, the relevance of relationships between nodes, the reliability of node summaries, and the alignment of confidence scores with the perceived strength and reliability of the connections. Each LLM independently rated the graph components on a scale of 1 to 10. The results showed an average accuracy score of 8.9/10 for node identification, 8.8/10 for relationship relevance, and 8.5/10 for the clarity and precision of node summaries. Confidence scores generally aligned well with the LLMs’s assessments, as illustrated in Tables 4 and 5, which highlight strong relationships across domains such as ophthalmology, cardiovascular treatments, and dermatology.
@@
The detailed evaluations in Tables 4 and 5 provide further insights into the graph’s performance across diverse medical domains. For instance, the accurate representation of relationships between beta-blockers like Labetalol and Propranolol or the integration of treatments such as Diltiazem and Nitroglycerin for cardiovascular care highlight the MKG’s capacity to support intricate clinical decision-making.
@@ Appendix C QA Samples with reasoning from MEDQA benchmark
Table 6 provides an example of how the model processes a clinical case question related to the management of acute coronary syndrome (ACS). The search items retrieved for possible answer choices (e.g., Nifedipine, Enoxaparin, Clopidogrel, Spironolactone, Propranolol) are accompanied by key content excerpts relevant to their roles in ACS treatment. Additionally, the reasoning pathways illustrate how the model synthesizes evidence-based knowledge to justify the selection of the correct answer (Clopidogrel), while also explaining why the alternative options are not suitable. Additional examples are also provided in Tables  7,  8, and  9

### v2

@@ 4.2 Performance Comparison
Table 1: Comparison of LLM models on the MEDQA Benchmark. Additional comparison with RAGs are provided in Table B
Table 1 presents a comprehensive comparison of state-of-the-art language models on the MEDQA benchmark. The results highlight the critical role of advanced reasoning strategies in achieving higher performance, such as CoT reasoning and the integration of search tools. While larger models like Med-Gemini and GPT-4 achieve the highest accuracy and F1 scores, their performance comes at the cost of significantly larger parameter sizes. These models exemplify the power of scaling combined with sophisticated reasoning and retrieval techniques.
Significantly, AMG-RAG, despite having just 8 billion parameters, attains an F1 score of 74.1% on the MEDQA benchmark, surpassing models like Meditron, which possess 70 billion parameters without needing any fine tuning. This highlights AMG-RAG’s exceptional efficiency and proficiency in utilizing CoT reasoning and external evidence retrieval. The model leverages tools such as PubMedSearch and WikiSearch to dynamically integrate domain-specific knowledge dynamically, thereby improving its ability to address medical questions. Examples of QA interactions, including detailed search items and reasoning for question samples, are provided in Appendix C. These examples are organized in Tables 7, 8, 9, and 10, drawn from the MEDQA benchmark.
@@
Overall, AMG-RAG’s results on MEDQA and MedMCQA benchmarks solidify its position as a highly efficient and effective model for medical QA. By leveraging reasoning, dynamically generated MKG, and external knowledge sources, AMG-RAG not only closes the gap with much larger models but also sets a new standard for performance among smaller-sized models.
Impact of Search Tools on MKG creation and CoT Reasoning on AMG-RAG Performance.
@@
Figure 3 and Table 3 demonstrate the effect of integrating different search tools for creating the MKG on the performance of the AMG-RAG system applied to the MEDQA benchmark. Incorporating these external retrieval capabilities significantly enhances both accuracy and F1 scores, as they allow the model to access relevant and up-to-date evidence critical for answering complex medical questions. Among the two search tools for creating the MKG, PubMed-MKG consistently outperforms Wiki-MKG, likely due to its focused, domain-specific content that aligns closely with the specialized nature of medical QA tasks.
In addition to the integration of the dynamical MKG, the reasoning module plays a pivotal role in performance. As highlighted in Figure 3, ablating either CoT or MKG integration causes a considerable degradation in accuracy and F1 score. This demonstrates that structured multi-hop reasoning and medical knowledge grounding through the MKG are indispensable for the system’s ability to deliver accurate and evidence-based answers."
@@
Figure 3: Confusion matrix for AMG-RAG with and without CoT and Knowledge Graph integration on the MEDQA dataset.
Comparison Against Traditional RAG Models.
Table 3 presents a comprehensive comparison of various RAG models evaluated on the MEDQA benchmark. This includes models with different retrieval mechanisms and model sizes, enabling a head-to-head evaluation of AMG-RAG with other state-of-the-art baselines such as Self-RAG (Asai et al., 2023), HyDE (Gao et al., 2022), GraphRAG (Edge et al., 2024), and MedRAG (Zhao et al., 2025). The results clearly show that AMG-RAG configured with the PubMed-MKG and an 8B LLM backbone achieves the highest accuracy of 73.92%, surpassing all competing models. Notably, ablation results indicate that removing search functionality or CoT reasoning significantly degrades accuracy (dropping to 67.16% and 66.69%, respectively), confirming the essential role of structured retrieval and reasoning components in complex question answering. Other baseline models such as Gemini-pro and PMC-LLaMA demonstrate weaker performance, further validating the efficacy of domain-aware retrieval and reasoning modules proposed in AMG-RAG. Importantly, the domain specificity and freshness of PubMedSearch provide a significant advantage in retrieving relevant knowledge that general-purpose search modules often fail to deliver.
Model Size Accuracy (%)
AMG-RAG PubMed-MKG-8B 73.92
Wiki-MKG-8B 70.62
No-MK-8B 67.16
No-MKG & CoT-8B 66.69
Self-RAG 8B 67.32
HyDE-8B 68.32
RAG Gemini-pro 64.5
70B 56.2
8B 64.3
GraphRAG Gemini-pro 65.1
70B 55.1
8B 64.8
MedRag 70B 49.57
13B 42.58
PMC-LLaMA 13B 44.38
Table 3: Comparison of MEDQA accuracy across various RAG models and retrieval strategies.
Comparison Against LLM Backbones.
In addition to evaluating different retrieval strategies, we assess how the choice of LLM backbone influences performance in Table 4. This comparison highlights that AMG-RAG built on GPT4o-mini with PubMed-MKG achieves the best performance (73.92%). In contrast, performance declines when switching to LLaMA 3.1 or Mixtral, even when using the same retrieval pipeline. These results reinforce the importance of synergy between the language model and the retrieval mechanism. Larger models do not necessarily guarantee higher accuracy—domain alignment and reasoning ability, such as that of GPT4o-mini, are crucial for success on high-stakes tasks like medical QA.
Model Config-Size Accuracy (%)
GPT4o-mini PubMed-MKG-8B 73.92
No-MKG & CoT-8B 66.69
LLaMA 3.1 PubMed-MKG-8B 66.5
No-MKG-8B 62.6
Mixtral PubMed-MKG-8$\times$7B 61.4
No-MKG-8$\times$7B 53.2
GPT 3.5 PubMed-MKG 65.2
No-MKG 58.4
Table 4: AMG-RAG performance across different LLM backbones on the MEDQA benchmark.
Improving QA in Rapidly Changing Medical Domains with AMG-RAG.
Figure 4 shows AMG-RAG’s superior performance in rapidly evolving subfields like Neurology and Genetics. This advantage stems from real-time PubMed integration during inference, combined with structured reasoning and knowledge graph grounding, enabling precise answers to complex medical questions with enhanced interpretability and trustworthiness.
@@ 5 Conclusion
We introduce AMG-RAG, an advanced QA system that dynamically constructs MKG while integrating sophisticated structured reasoning for medical QA. The system demonstrates significant improvements in accuracy and reasoning capabilities, particularly for medical question-answering tasks, outperforming other approaches of similar model size or 10 to 100 times larger.
@@ 6 Limitations
Despite AMG-RAG’s advancements, our approach has certain limitations. Firstly, it relies on external search tools which introduce latency during the creation of the MKG. However, this occurs only once, when the MKG is built from scratch for the first time. Additionally, while the system performs exceptionally well in medical domains, its applicability to non-medical tasks remains unexplored.
@@ Appendix B Evaluating the Accuracy and Robustness of the Medical Knowledge Graph
To assess accuracy and robustness, a two-phase evaluation process was employed. In the first phase, a group of expert LLMs specialists in medical domains reviewed a subset of the MKG, including dynamically generated nodes, relationships, confidence scores, and summaries for various medical queries. They evaluated the accuracy of medical terms and concepts, the relevance of relationships between nodes, the reliability of node summaries, and the alignment of confidence scores with the perceived strength and reliability of the connections. Each LLM independently rated the graph components on a scale of 1 to 10. The results showed an average accuracy score of 8.9/10 for node identification, 8.8/10 for relationship relevance, and 8.5/10 for the clarity and precision of node summaries. Confidence scores generally aligned well with the LLMs’s assessments, as illustrated in Tables 5 and 6, which highlight strong relationships across domains such as ophthalmology, cardiovascular treatments, and dermatology.
@@
The detailed evaluations in Tables 5 and 6 provide further insights into the graph’s performance across diverse medical domains. For instance, the accurate representation of relationships between beta-blockers like Labetalol and Propranolol or the integration of treatments such as Diltiazem and Nitroglycerin for cardiovascular care highlight the MKG’s capacity to support intricate clinical decision-making.
@@ Appendix C QA Samples with reasoning from MEDQA benchmark
Table 7 provides an example of how the model processes a clinical case question related to the management of acute coronary syndrome (ACS). The search items retrieved for possible answer choices (e.g., Nifedipine, Enoxaparin, Clopidogrel, Spironolactone, Propranolol) are accompanied by key content excerpts relevant to their roles in ACS treatment. Additionally, the reasoning pathways illustrate how the model synthesizes evidence-based knowledge to justify the selection of the correct answer (Clopidogrel), while also explaining why the alternative options are not suitable. Additional examples are also provided in Tables  8,  9, and  10
@@ E.1 Neo4j
Does Neo4j-Based Storage scale well?
Neo4j’s scalability for our medical knowledge graph storage is strategically robust, offering several key advantages for large-scale, relationship-intensive medical data. Its graph-based architecture is particularly well-suited for handling highly interconnected medical knowledge networks, supporting horizontal scaling that enables efficient performance even as the knowledge base grows. The cloud-based accessibility further enhances the framework’s flexibility, allowing seamless knowledge sharing and distributed access without local storage constraints.
How large were the knowledge graphs?
Our automatically constructed medical knowledge graphs demonstrate significant complexity and depth, comprising approximately 76,681 nodes and 354,299 edges. These nodes encompass a comprehensive range of medical entities including diseases, symptoms, treatments, drugs, anatomical structures, and clinical findings, all interconnected through semantically meaningful, typed relationships. This substantial scale not only reflects the intricate nature of medical knowledge but also enables more nuanced, multi-hop reasoning capabilities across diverse medical queries. The graph’s architecture allows for dynamic expansion and refinement, ensuring that the knowledge representation remains both comprehensive and adaptable to emerging medical research and understanding.
Appendix F Additional Results
Example 1 Question: A 29-year-old man presents with infertility. He has been trying to conceive for over 2 years. His wife has no fertility issues. Exam shows tall stature, long limbs, sparse body hair, gynecomastia, and small testes. Labs reveal elevated FSH and a high estradiol:testosterone ratio. Cytogenetic analysis indicates a chromosomal abnormality. If inherited from the father, during which stage of spermatogenesis did this error most likely occur?
Choices: A: Primary spermatocyte, B: Secondary spermatocyte, C: Spermatid, D: Spermatogonium, E: Spermatozoon
Answer: A (Primary spermatocyte)
Reasoning: This corresponds to an error in meiosis I during the father’s spermatogenesis, consistent with Klinefelter syndrome due to paternal nondisjunction.
Retrieved Papers: 1) Black et al., *The Genetic Landscape of Male Factor Infertility*, Uro, 2025. 2) Niyaz et al., *Chromosome Disorders in Sperm Anomalies*, 2025. 3) Leslie et al., *MNS1 variant and Male Infertility*, EJHG, 2020.
Example 2 Question: A 23-year-old woman is referred for genetic counseling after her brother is diagnosed with hereditary hemochromatosis. She is asymptomatic and her labs are normal. Which gene mutation is most consistent with hereditary hemochromatosis?
Choices: A: BCR-ABL, B: BRCA, C: FA, D: HFE, E: WAS
Answer: D (HFE gene)
Reasoning: Most hereditary hemochromatosis cases in Northern European populations are caused by HFE mutations (C282Y, H63D). Even asymptomatic individuals with normal iron studies should be screened if they have an affected first-degree relative.
Retrieved Papers: 1) Delatycki & Allen, *Population Screening for HH*, Genes, 2024. 2) Lou et al., *Utility of Iron Indices in HH Genotyping*, Clin. Biochem., 2025. 3) Lucas et al., *HFE Genotypes and Outcomes*, BMJ Open, 2024.
Table 11: Examples of AMG-RAG-generated answers with structured reasoning and citation-based grounding for clinical QA.